# Gradient Descent with Large Step Sizes: Chaos and Fractal Convergence Region

### v1

@@ E.1.2 Proof of Theorem 1
Also, when $Q(z,w)>8$ or $|\mu|>1$, almost all initializations do not converge. This gives the critical step size (3).
@@
The last equation is a quadratic form, and the eigenvalues of the coefficient matrix are $\eta\pm\frac{1}{2}\eta^{2}|y|$, each with multiplicity $d$. Therefore, when $\eta|y|<1$, the quadratic form is positive definite and defines a smooth ellipsoid of dimension 2d12d-1. Notice that, $\left\{u=\pm v\right\}$ is the union of two linear subspace with dimension $d$. Therefore, since $2d-1\geq 1$ and an ellipsoid is curved everywhere, there always exits a point $\bar{\theta}\in W\cap\partial\mathcal{D}^{\prime}_{\eta}\setminus\left\{u=\pm v\right\}$ and a neighborhood $\bar{W}$ of $\bar{\theta}$ such that $\bar{W}\subset W$ and $\bar{W}\cap\left\{u=\pm v\right\}=\varnothing$. The Jacobian of $T$ is full rank at all points in $\bar{W}$, so by constant rank theorem, $T(\bar{W})$ is an open set. Meanwhile, $T(\bar{\theta})\in T(\partial\mathcal{D}_{\eta}^{\prime})$. Hence, under the conjugacy (7), we have $T(\bar{W})\cap\left\{w=\mu z+4\right\}\neq\varnothing$. According to Proposition 20, there exists $(z^{\prime},w^{\prime}),(z^{\prime\prime},w^{\prime\prime})\in T(\bar{W})$ such that $F^{N}(z^{\prime},w^{\prime})$ converges to $(0,w^{*})$ with any $w^{*}\in[2|\mu|,\infty)$ and $F^{N}(z^{\prime\prime},w^{\prime\prime})$ converges to $(-\mu,0)$, as $N$ tends to infinity. Therefore, there exists $\theta^{\prime},\theta^{\prime\prime}\in W$ such that $\mathrm{GD}_{\eta}^{N}(\theta^{\prime})$ converges to a global minimizer with squared norm in $2|y|,\infty)$ and $\mathrm{GD}_{\eta}^{N}(\theta^{\prime\prime})$ converges to $(\boldsymbol{0},\boldsymbol{0})$. Notice that
@@
Hence, at any global minimizer, the imbalance is lower bounded by the squared norm. Hence, by what we have shown, arbitrarily large imbalance can be also attained by initialization in $\bar{W}$.
Finally, the result of topological entropy and the existence of periodic orbit of any period directly come from Proposition 18. This completes the proof. ∎
Lastly, we present a basic property for the unregularized scalar factorization problem: the sharpness coincides with the squared norm at the set of global minimizers. This property has been proved by Wang et al., (2022).
@@ E.2.1 Preliminary results
Proposition 22 (Small step size).
Assume $0<\nu<1-|\mu|$. For almost all $(z,w)\in\Omega$, if $Q(z,w)<8-4\nu$, we have that $F^{N}(z,w)$ converges to the global minimizer. If $Q(z,w)<4-4\nu$, we have that, for any $(u,v)$ that satisfies $T(u,v)=(z,w)$, $\mathrm{GD}_{\eta}^{N}(u,v)$ converges to $p^{-}(u,v)$, as defined in Theorem 3.
@@
By Lemma 12, we have that $Q(F^{N}(z,w))$ monotonically decreases and converges to $\left\{\Delta Q=0\right\}=\left\{Q=4|\mu|\right\}$. As shown in proving Lemma 11, the $w$-update on $\left\{Q=4|\mu|\right\}$ is given by
@@
By differentiation, when $|\mu|>\nu$, $\kappa(w)$ has two fixed points $w=0$ and $w=2(|\mu|-\nu)$ on $0,4|\mu|]$. In particular, $w=0$ is repelling and $w=2(|\mu|-\nu)$ is globally attracting, i.e., attracting all orbits on $0,4|\mu|]$ except $w=0$. When $|\mu|\leq\nu$, $\kappa(w)$ has only one fixed point $w=0$ which is globally attracting. In either case, the attracting fixed points is the global minimizer and hence for almost all initializations, the orbit converges to the minimizer. The measure-zero event occurs if $|\mu|>\nu\geq 0$ and the initializations lies on the measure-zero set (it has measure zero since the Jacobian of $F$ has full rank almost everywhere):
@@
for some neighborhood $O$ of $(-\mu,0)$. Note, this set is the basin of attraction of the saddle $(-\mu,0)$
Without loss of generality, assume $y\geq 0$. Consider $p=(u+v)/\sqrt{2}$ and $q=(u-v)/\sqrt{2}$. Note that when $y>\lambda$, the set of global minimizers is given by
@@
Also, notice that
$\sqrt{2}\cdot p^{\prime}=u^{\prime}+v^{\prime}=u-\eta zv-\eta\lambda u+v-\eta zu-\eta\lambda v=(1-\eta z-\eta\lambda)(u+v)=(1-\eta z-\eta\lambda)\sqrt{2}\cdot p.$
@@
Similarly,
$q_{k}=q_{0}\Pi_{j=0}^{k-1}(1+z_{j}-\nu).$
Therefore, the converged point is either $(\sqrt{2(y-\lambda)}\frac{p_{0}}{\|p_{0}\|},0)$ or $(-\sqrt{2(y-\lambda)}\frac{p_{0}}{\|p_{0}\|},0)$. Since the global minimizer is given by $\left\{q=0,\|p\|^{2}=2(y-\lambda)\right\}$, the former is the minimal distance solution and the latter is the maximal distance solution. Note the change of coordinates $p=(u+v)/\sqrt{2}$ and $q=(u-v)/\sqrt{2}$ is given by an orthogonal transformation which preserves distance. Therefore the same statement holds in the $uv$-coordinate.
@@
Note the line intersects with $\partial\Omega$ at $(1-\nu,2(1-\nu+\mu))$ and $(\nu-1,-2(\nu-1+\mu))$. Therefore we have that for all initializations that satisfy $Q(z,w)<4-4\nu$, we have $|z|<1-\nu$. Meanwhile, since $4-4\nu<8-4\nu$, we have that the $\{Q<4-4\nu\}$ is forward invariant by Lemma 12, i.e., $|z|<1-\nu$ holds on the entire orbit. Therefore, when $Q(z,w)<4-4\nu$, we have that $1\pm z_{j}-\nu>0$ for all $j\geq 0$. It follows that, by (17), the converged minimizer in $pq$-coordinate has to be $(\sqrt{2(y-\lambda)}\frac{p_{0}}{\|p_{0}\|}$, which is the minimal distance solution. This completes the proof. ∎
@@
We use $\mathcal{D}$ for $\mathcal{D}_{\eta}^{\prime\prime}$ for notation simplicity. First, we prove that $T(\partial\mathcal{D}_{\eta}^{\prime\prime})=\partial T(\mathcal{D_{\eta}^{\prime\prime}})$. Notice that, by Proposition 8, $T^{-1}(T(\mathcal{D}))=\mathcal{D}$, i.e., $\mathcal{D}$ is saturated with respect to $T$. Since $T$ is continuous, we have that $\partial T^{-1}(A)\subset T^{-1}(\partial A)$ for any set $A$. Hence, $\partial T^{-1}(T(\mathcal{D}))=\partial\mathcal{D}\subset T^{-1}(\partial T(\mathcal{D}))$. It follows that, $T(\partial\mathcal{D})\subset\partial T(\mathcal{D})$. By Proposition 22,
$\Gamma=\left\{(u,v)\colon Q(T(u,v))=\left\lVert u\right\rVert^{2}+\left\lVert v\right\rVert^{2}+\sqrt{(\left\lVert u\right\rVert^{2}+\left\lVert v\right\rVert^{2})^{2}-16y(u^{\top}v-y)}<8-4\lambda\right\}$
is an open set and all points in this set converge to the global minimizer or the saddle. Hence, $\mathcal{D}$ is an open set. To see this, if $x\in\mathcal{D}$ then we have for sufficiently large $N$, $\mathrm{GD}^{N}_{\eta}(x)\in\Gamma$. By the continuity of $\mathrm{GD}^{N}_{\eta}$, we have $x$ is in an interior point of $\mathcal{D}$. Now consider any point $y\in\partial T(\mathcal{D})$. For the sake of contradiction, assume $y\notin T(\partial\mathcal{D})$. Then we have $T^{-1}(y)\cap\partial\mathcal{D}=\varnothing$. Note $T^{-1}(y)$ is a compact set by the properness of $T$. Since $\mathcal{D}$ is open, there exists an open neighborhood $O$ of $T^{-1}(y)$ such that $O\subset\mathcal{D}$ or $O\subset(\mathcal{D}^{c})^{o}$. In either case, we have $y\in T(O)$ and $T(O)\cap T(\partial\mathcal{D})=\varnothing$, which yields a contradiction. Hence $y\in T(\partial\mathcal{D})$ and $\partial T(\mathcal{D})\subset T(\partial\mathcal{D})$. It follows that $\partial T(\mathcal{D})=T(\partial\mathcal{D})$.
Next, we prove that, $F(\partial T(\mathcal{D}))=\partial T(\mathcal{D})$. Now let $A=T(\mathcal{D})$ for simplicity. First, note that $A$ satisfies $F(A)\subset A$, $F^{-1}(A)\subset A$ by the definition of $A$. Consider any point $x$ such that $T(x)\in\partial A$. Then any small enough neighborhood $O$ of $x$ is mapped a neighborhood of $T(x)$, which contains a point $y\in A$ and a point $z\in A^{c}$. Since $F^{-1}(A)\subset A$, $F^{-1}(y)\in O\cap A$. Since $F(A)\subset A$, $F^{-1}(z)\in A^{c}$. Therefore, we have that $F^{-1}(\partial A)\subset\partial A$. Meanwhile, since $F$ is surjective by Proposition 13, we have $F\circ F^{-1}(\partial A)=\partial A$. It follows that
@@
On the other side, for any $y\in\partial A$, since $y\in\mathrm{cl}(A)$ and $F$ is continuous, we have that $T(y)\in\mathrm{cl}(F(A))=\mathrm{cl}(A)$. Since $F^{-1}(A)\subset A$, $T(y)\notin A$. Hence, $T(y)\in\partial A$. This means that $F(\partial A)\subset\partial A$. Therefore, $F(\partial A)=\partial A$.
@@
where $G_{i}$’s are homeomorphisms. As shown in Proposition 13, $G_{i}(\Omega^{o})\cap G_{j}(\Omega^{o})$ is empty whenever $i\neq j$. Therefore, $\partial T(\mathcal{D}_{\eta}^{\prime\prime})$ is self-similar with degree three. This completes the proof. ∎
In the following, we show that $\mathcal{D}_{\eta}$ is equal almost everywhere to a set with unbounded interior.
@@
$|z^{\prime}|$ $=|z|\cdot|z^{2}-w\alpha+\alpha^{2}|$
@@
$w^{\prime}$ $\leq w(z^{2}+\alpha^{2})+4z^{2}\alpha$
@@
Hence,
@@
$\Leftarrow$ $1+b(w+4\alpha)a^{2}\exp(-2bw)>a^{2}\exp((\alpha^{2}-3)bw)+\exp((\alpha^{2}-1)bw)(\alpha^{2}+\alpha w).$
@@
$p(w)=1+b(w+4\alpha)a^{2}\exp(-2bw),\ q(w)=a^{2}\exp((\alpha^{2}-3)bw)+\exp((\alpha^{2}-1)bw)(\alpha^{2}+\alpha w).$
@@
$p^{\prime}(w)$ $=a^{2}b\exp(-2bw)\Big(1-2b(w+4\alpha)\Big)$
$\propto-2bw+1-8\alpha b.$
Therefore $p(w)$ increases from $(-\infty,w_{0})$ for some $w_{0}\in\mathbb{R}$ and decreases on $w_{0},+\infty)$. Since $\lim_{w\to+\infty}p(w)=1$, we have that
$\min_{w\in[0,+\infty)}p(w)=\min\left\{p(0),1\right\}=\min\{1+4a^{2}b\alpha,1\}=1.$
@@
$\propto a^{2}(\alpha^{2}-3)b\exp(-2bw)+(\alpha^{2}-1)b\alpha w+(\alpha^{2}-1)b\alpha^{2}+\alpha.$
[…]
@@
Since $0<\alpha<1$, we can always select $a$ small enough such that $q(0)<1$. Under such selection we have that $p(w)>q(w)$ holds for all $w\geq 0$ and hence (18) holds. Therefore, $\left\{|z|<a\exp(-bw)\right\}$ is forward invariant. Due to the exponential decay, we can always select $a$ small enough and $b$ large enough such that $\left\{|z|<a\exp(-bw)\right\}\subset\left\{Q(z,w)>Q(F(z,w))\right\}$, where the latter set is given in Lemma 12. Therefore, in this exponential cone, $Q$ monotonically decays, and the orbit converges to the minimizer almost surely. This completes the proof. ∎
@@
Let $(z^{\prime},w^{\prime})=F(z,w)$ and $m_{k}=G^{k}_{0}(m_{0})$. Consider the function $E(z,w)=w+2(z+\mu)$. We have
$E(F(z,w))-E(z,w)=$ $w^{\prime}+2(z^{\prime}+\mu)-w-2(z+\mu)$
@@
Note for $(z,w)\in\Omega$, $w+2(z+\mu)\geq 0$. When $w>2(z+\mu)$, we have that
$E(F(z,w))-E(z,w)>0\Leftarrow z<-\nu.$ (19)
We have that $m_{k}\in\mathrm{cl}(A_{0})$ for $k\geq 1$. Hence, the $z$-coordinate of $m_{k}$ is smaller than $\nu-1$ for all $k\geq 1$. Since $\nu<1/2$, $\nu-1<-\nu$. Hence, $m_{k}\in\left\{E(F(z,w))>E(z,w)\right\}$ for all $k\geq 1$. This implies that $E(m_{k})$ monotonically decreases. Since $E$ has lower bound 0 on $\Omega$, $E(m_{k})$ converges to some finite value $E^{*}$. For contradiction, assume $m_{k}$ is unbounded. Note for any $m_{1}$, we have the set
@@
is bounded for any $M>0$. Hence, we have the $z$-coordinate of $m_{k}$ tends to negative infinity. Note that for sufficiently small $z$ and $(z,w)\in\Omega$, we have

### v2

@@ E.1.2 Proof of Theorem 1
Also, by Proposition 15, when $Q(z,w)>8$ or $|\mu|>1$, almost all initializations do not converge. This gives the critical step size (3).
@@
The last equation is a quadratic form, and the eigenvalues of the coefficient matrix are $\eta\pm\frac{1}{2}\eta^{2}|y|$, each with multiplicity $d$. Therefore, when $\eta|y|<1$, the quadratic form is positive definite and defines a smooth ellipsoid of dimension 2d12d-1. Notice that, $\left\{u=\pm v\right\}$ is the union of two linear subspace with dimension $d$. Therefore, since $2d-1\geq d$ and an ellipsoid is curved everywhere, there always exits a point $\bar{\theta}\in W\cap\partial\mathcal{D}^{\prime}_{\eta}\setminus\left\{u=\pm v\right\}$ and a neighborhood $\bar{W}$ of $\bar{\theta}$ such that $\bar{W}\subset W$ and $\bar{W}\cap\left\{u=\pm v\right\}=\varnothing$. The Jacobian of $T$ is full rank at all points in $\bar{W}$, so by constant rank theorem, $T(\bar{W})$ is an open set. Meanwhile, $T(\bar{\theta})\in T(\partial\mathcal{D}_{\eta}^{\prime})$. Hence, under the conjugacy (7), we have $T(\bar{W})\cap\left\{w=\mu z+4\right\}\neq\varnothing$. According to Proposition 20, there exists $(z^{\prime},w^{\prime}),(z^{\prime\prime},w^{\prime\prime})\in T(\bar{W})$ such that $F^{N}(z^{\prime},w^{\prime})$ converges to $(0,w^{*})$ with any $w^{*}\in[2|\mu|,\infty)$ and $F^{N}(z^{\prime\prime},w^{\prime\prime})$ converges to $(-\mu,0)$, as $N$ tends to infinity. Therefore, there exists $\theta^{\prime},\theta^{\prime\prime}\in W$ such that $\mathrm{GD}_{\eta}^{N}(\theta^{\prime})$ converges to a global minimizer with squared norm in $2|y|,\infty)$ and $\mathrm{GD}_{\eta}^{N}(\theta^{\prime\prime})$ converges to $(\boldsymbol{0},\boldsymbol{0})$. Notice that
@@
Hence, at any global minimizer, the imbalance is lower bounded by the squared norm. Hence, by what we have shown, arbitrarily large imbalance can be also attained by initializations in $\bar{W}$.
Finally, the lower bound for the topological entropy and the existence of periodic orbit of any period directly come from Proposition 18. This completes the proof. ∎
Lastly, we present a basic property for the unregularized scalar factorization problem: the sharpness coincides with the squared norm at the set of global minimizers. This has been proved by Wang et al., (2022). Below we represent their results.
@@ E.2.1 Preliminary results
Proposition 22.
Assume $0<\nu<1-|\mu|$. For almost all $(z,w)\in\Omega$, if $Q(z,w)<8-4\nu$, we have that $F^{N}(z,w)$ converges to $T(\mathcal{M})$ as $N\to\infty$. If $Q(z,w)<4-4\nu$, for any $(u,v)$ that satisfies $T(u,v)=(z,w)$, $\mathrm{GD}_{\eta}^{N}(u,v)$ converges to $p^{-}(u,v)$, as defined in Theorem 3, as $N\to\infty$.
@@
By Lemma 12, we have that $Q(F^{N}(z,w))$ monotonically decreases. Since $Q(F^{N}(z,w))$ is non-negative, it converges to some finite value. It follows that $F^{N}(z,w)$ converges to $\left\{\Delta Q=0\right\}\cap\left\{Q<8-4\nu\right\}$, which is equal to $\left\{Q=4|\mu|\right\}$ according to Lemma 12. Note $\left\{Q=4|\mu|\right\}=\left\{w=2\mathrm{sgn}(\mu)(z+\mu),w\leq 4|\mu|\right\}$ by Lemma 10. The $w$-update under $F$ on $\left\{Q=4|\mu|\right\}$ is given by
@@
When $|\mu|>\nu$, $\kappa(w)$ has two fixed points on $0,4|\mu|]$: $w=0$ and $w=2(|\mu|-\nu)$. It is straight forward to obtain that $w=2(|\mu|-\nu)$ attracts all orbits on $0,4|\mu|]$ except the one with initial value $w=0$. Note $w=0$ corresponds to the saddle $(-\mu,0)$, where the Jacobian of $F$ has eigenvalues: $(1+\mu-\nu)^{2}$ and $(-1+\mu+\nu)^{2}$. Hence, $(-\mu,0)$ is a hyperbolic fixed point of $F$. Note also that, $w=2(|\mu|-\nu)$ corresponds to $(-\mathrm{sgn}(\mu)\nu,2(|\mu|-\nu))$, which is the only element of $T(\mathcal{M})$. The Jacobian of $F$ at this point has eigenvalues: $(-1+2\nu)^{2},$ and $1-2|\mu|+2\nu$. Hence, $(-\mathrm{sgn}(\mu)\nu,2(|\mu|-\nu))$ attracts nearby orbits. Then with the omega-limit set arguments similar to those in the proof of Proposition 15, we have that, for all initializations that do not lie in the basin of attraction of $(-\mu,0)$:
@@
where $O$ is a neighborhood of $(-\mu,0)$, the orbit converges to $T(\mathcal{M})$. The above basin has measure-zero since the Jacobian of $F$ has full rank almost everywhere. Therefore, for almost all initializations with $Q(z,w)<8-4\nu$, the orbit converges to $T(\mathcal{M})$. When $|\mu|\leq\nu$, $\kappa(w)$ has only one fixed point $w=0$, which attracts all orbits $0,4|\mu|]$. Note in this case, $(-\mu,0)$ attracts nearby orbits, since both eigenvalues of $JF$ have norm smaller than one. Then, with the omega-limit set arguments, we have that for all initializations with $Q(z,w)<8-4\nu$, the orbit converges to $T(\mathcal{M})$.
Next, we identify the converged point under $\mathrm{GD}_{\eta}$ when $Q<4-4\nu$. Without loss of generality, assume $y\geq 0$. Let $(z_{k},w_{k})_{k\geq 0}$ be the orbit and $T(u_{k},v_{k})=(z_{k},w_{k})$. Consider new coordinates $p=(u+v)/\sqrt{2}$ and $q=(u-v)/\sqrt{2}$. Note that when $y>\lambda$, the set of global minimizers is given by
@@
Notice that
$\sqrt{2}\cdot p_{k+1}=u_{k+1}+v_{k+1}$ $=u_{k}-\eta z_{k}v_{k}-\eta\lambda u_{k}+v_{k}-\eta z_{k}u_{k}-\eta\lambda v_{k}$
$=(1-\eta z_{k}-\eta\lambda)(u_{k}+v_{k})=(1-\eta z_{k}-\eta\lambda)\sqrt{2}\cdot p_{k}.$
@@
Therefore, the converged point is either $(\sqrt{2(y-\lambda)}\frac{p_{0}}{\|p_{0}\|},0)$ or $(-\sqrt{2(y-\lambda)}\frac{p_{0}}{\|p_{0}\|},0)$. Since the global minimizer is given by $\left\{q=0,\|p\|^{2}=2(y-\lambda)\right\}$, the former point is the minimal distance solution and the latter point is the maximal distance solution, under the $pq$-coordinates. Note the change of coordinates $p=(u+v)/\sqrt{2}$ and $q=(u-v)/\sqrt{2}$ is given by an orthogonal transformation which preserves distance. Therefore the same statement holds in the $uv$-coordinates.
@@
The above line intersects with $\partial\Omega$ at $(1-\nu,2(1-\nu+\mu))$ and $(\nu-1,-2(\nu-1+\mu))$. Therefore, for all initializations that satisfy $Q(z,w)<4-4\nu$, we have $|z|<1-\nu$. Meanwhile, since $4-4\nu<8-4\nu$, we have that the $\{Q<4-4\nu\}$ is forward invariant by Lemma 12, i.e., $|z|<1-\nu$ holds on the entire orbit. This implies that, when $Q(z_{0},w_{0})<4-4\nu$, we have that $1\pm z_{j}-\nu>0$ for all $j\geq 0$. By (22), the converged minimizer in $pq$-coordinate has to be $(\sqrt{2(y-\lambda)}\frac{p_{0}}{\|p_{0}\|},0)$, which is the minimal distance solution. This completes the proof. ∎
@@
We use $\mathcal{D}$ for $\mathcal{D}_{\eta}^{\prime\prime}$ for notation simplicity. First, we prove that $T(\partial\mathcal{D}_{\eta}^{\prime\prime})=\partial T(\mathcal{D_{\eta}^{\prime\prime}})$. Notice that, by Proposition 8, orbit under $\mathrm{GD}_{\eta}$ converges to $\mathcal{M}$ or to the saddle if and only if the corresponding orbit under $F$ converges to $T(\mathcal{M})$ or to $(-\mu,0)$, respectively. This implies that $T^{-1}(T(\mathcal{D}))=\mathcal{D}$ and $T(\mathcal{D}^{c})=T(\mathcal{D})^{c}$. Since $T$ is continuous, we have that $\partial T^{-1}(A)\subset T^{-1}(\partial A)$ for any set $A$. Hence, $\partial T^{-1}(T(\mathcal{D}))=\partial\mathcal{D}\subset T^{-1}(\partial T(\mathcal{D}))$. It follows that, $T(\partial\mathcal{D})\subset\partial T(\mathcal{D})$.
For the other direction, consider any point $y\in\partial T(\mathcal{D})$ and assume $y\notin T(\partial\mathcal{D})$ for contradiction. Then we have $T^{-1}(y)\cap\partial\mathcal{D}=\varnothing$. Note $T^{-1}(y)$ is a connected set by Proposition 7. Thus, there exists an open neighborhood $O$ of $T^{-1}(y)$ such that $O\subset\mathcal{D}^{o}$ or $O\subset(\mathcal{D}^{c})^{o}$. Hence, $T(O)\subset T(\mathcal{D})$ or $T(O)\subset T(\mathcal{D}^{c})=T(\mathcal{D})^{c}$, which contradicts the fact that $y\in\partial T(\mathcal{D})$. Hence $y\in T(\partial\mathcal{D})$ and $\partial T(\mathcal{D})\subset T(\partial\mathcal{D})$. It follows that $\partial T(\mathcal{D})=T(\partial\mathcal{D})$.
Next, we prove that, $F(\partial T(\mathcal{D}))=\partial T(\mathcal{D})$. Now let $A=T(\mathcal{D})$ for simplicity. Note that $A$ satisfies $F(A)\subset A$, $F^{-1}(A)\subset A$, since $A=T(\mathcal{D})$ contains all initializations that converge to $T(\mathcal{M})$ and to $(-\mu,0)$. Consider any point $x\in F^{-1}(\partial A)$. Then a small enough neighborhood $O$ of $x$ is mapped to a neighborhood of $F(x)$, which contains a point $y\in A$ and a point $z\in A^{c}$. Since $F^{-1}(A)\subset A$, $F^{-1}(y)\in O\cap A$. Since $F(A)\subset A$, $F^{-1}(z)\in O\cap A^{c}$. Therefore, $x\in\partial A$ and $F^{-1}(\partial A)\subset\partial A$. Meanwhile, since $F$ is surjective by Proposition 13, we have $F\circ F^{-1}(\partial A)=\partial A$. It follows that
@@
For the other direction, note that for any $y\in\partial A$, since $y\in\mathrm{cl}(A)$ and $F$ is continuous, we have that $F(y)\in\mathrm{cl}(F(A))\subset\mathrm{cl}(A)$. Consider the set $\left\{Q(z,w)<8-4\nu\right\}$. By Proposition 22, $\left\{Q(z,w)<8-4\nu\right\}$ is an open neighborhood of $T(\mathcal{M})\cup\{(\mu,0)\}$ and all orbits in this neighborhood converge to $T(\mathcal{M})\cup\{(\mu,0)\}$. Now consider any $x\in A$. Since $F^{N}(x)$ converges to $T(\mathcal{M})$ or to $(\mu,0)$, $F^{N}(x)$ must visit $\Gamma$ at some $N$. Since $\Gamma$ is open and $F^{N}$ is continuous, $x$ is an interior point of $A$. Thus, $A$ is an open set. Therefore, $y\in\mathrm{cl}(A)\setminus A^{o}=\mathrm{cl}(A)\setminus A$. Since $F^{-1}(A)\subset A$, $F(y)\notin A$. Taken together, we have that $F(y)\in\mathrm{cl}(A)\setminus A=\partial A$ and $F(\partial A)\subset\partial A$. Therefore, $F(\partial A)=\partial A$.
@@
where $G_{i}$’s are homeomorphisms. As shown in Proposition 13, $G_{i}(\Omega^{o})\cap G_{j}(\Omega^{o})$ is empty whenever $i\neq j$. Therefore, $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$ is self-similar with degree three. This completes the proof. ∎
In the following, we show that $\mathcal{D}_{\eta}$ has an unbounded interior, up to a measure-zero set.
@@
$|z^{\prime}|$ $=|z^{3}+(\alpha^{2}-\alpha w)z|$
$=|z|^{3}+(\alpha^{2}+\alpha w)|z|$
@@
$w^{\prime}$ $=w(z^{2}+\alpha^{2})-4z^{2}\alpha$
@@
Note $a\exp(-bw^{\prime})$ decreases as $w^{\prime}$ increases. Then,
@@
$\Leftarrow$ $1-b(w+4\alpha)a^{2}\exp(-2bw)>a^{2}\exp((\alpha^{2}-3)bw)+\exp((\alpha^{2}-1)bw)(\alpha^{2}+\alpha w).$
@@
$p(w)=1-b(w+4\alpha)a^{2}\exp(-2bw),\quad q(w)=a^{2}\exp((\alpha^{2}-3)bw)+\exp((\alpha^{2}-1)bw)(\alpha^{2}+\alpha w).$
@@
$p^{\prime}(w)$ $=-a^{2}b\exp(-2bw)\Big(1-2b(w+4\alpha)\Big)$
$\propto 2bw-1+8\alpha b.$
Therefore $p(w)$ decreases from $(-\infty,w_{0})$ and increases on $w_{0},+\infty)$, where $w_{0}=(1-8\alpha b)/(2b)$. Let $b>1/(8\alpha)$. Then we have that
$\min_{w\in[0,+\infty)}p(w)=p(0)=1-4\alpha a^{2}b.$
@@
$\propto a^{2}(\alpha^{2}-3)b\exp(-2bw)+(\alpha^{2}-1)b\alpha w+(\alpha^{2}-1)b\alpha^{2}+\alpha$
$\propto a^{2}(\alpha^{2}-3)b\exp(-2bw)+\Big((\alpha^{2}-1)\alpha w+(\alpha^{2}-1)\alpha^{2}\Big)b+\alpha.$
Note, that $a^{2}(\alpha^{2}-3)b\exp(-2bw)<0$ always holds. Also, since $(\alpha^{2}-1)\alpha w+(\alpha^{2}-1)\alpha^{2}\leq 0$ when $w\geq 0$, we can choose $b$ sufficiently large so that $q^{\prime}(w)<0$ holds on $0,+\infty)$. Then
@@
Now fix $b$. Note $1-4\alpha a^{2}b\to 1$ and $a^{2}+\alpha^{2}\to\alpha^{2}$ as $a\to 0$. We can always select $a$ small enough such that $q(0)<p(0)$. Then we have that $p(w)>q(w)$ holds for all $w\geq 0$ and hence (23) holds. Therefore, the set $\left\{|z|<a\exp(-bw)\right\}$ is forward invariant under $F$. Due to the exponential decay, we can always select $a$ small enough and $b$ large enough such that $\left\{|z|<a\exp(-bw)\right\}\subset\left\{Q(z,w)>Q(F(z,w))\right\}$, where the latter set is given in Lemma 12. Therefore, in this exponential cone, $Q$ decreases monotonically. Then using similar arguments to those in Proposition 22, we have that all almost all initializations in this cone converge to the minimizer. This completes the proof. ∎
@@
We prove the case $\mu\geq 0$. Note that the case of $\mu<0$ can be proved via analogous procedures. Let $(z^{\prime},w^{\prime})=F(z,w)$ and $m_{k}=G^{k}_{0}(m_{0})$ for $k\geq 1$. Consider the function $E(z,w)=w+2(z+\mu)$. We have
$E(F(z,w))-E(z,w)=$ $w^{\prime}+2(z^{\prime}+\mu)-w-2(z+\mu)$ (24)
@@
Note for $(z,w)\in\Omega$, $w+2(z+\mu)\geq 0$. Also, $2-\nu>0>-\nu$. Then for $w>2(z+\mu)$, we have that
$E(F(z,w))-E(z,w)>0\Leftarrow z<-\nu.$
We have that $m_{k}\in\mathrm{cl}(A_{0})$ for $k\geq 1$. Hence, the $z$-coordinate of $m_{k}$ is smaller than $\nu-1$ for all $k\geq 1$. Since $\nu<1/2$, $\nu-1<-\nu$. Hence, $m_{k}\in\left\{E(F(z,w))>E(z,w)\right\}$ for all $k\geq 1$. Note $(m_{k})_{k\geq 0}$ is a backward orbit. Thus, $E(m_{k})$ monotonically decreases. Since $E$ is lower bounded by 0 on $\Omega$, $E(m_{k})$ converges to some finite value $E^{*}$. For contradiction, assume $m_{k}$ is unbounded. Note that, given arbitrary $m_{1}\in\mathrm{cl}(A_{0})$, the set
@@
is bounded for any $M>0$. Hence, we have the $z$-coordinate of $m_{k}$ tends to negative infinity. Note that for sufficiently small negative $z$ and $(z,w)\in\Omega$, we have